How to solve systems by substitution
y=x-2
y=4x+1

Answer by checkley79(3341)

y=x-2
y=4x+1
x-2=4x+1
x-4x=1+2
-3x=3
x=3/-3
x=-1 ans.
y=-1-2
y=-3 ans.
Proof:
-3=4*-1+1
-3=-4+1
-3=-3